OpenStudy (anonymous):

the mass of 4Be7 is 1.1652*10, what is the defect of the beryllium nucleus?

OpenStudy (anonymous):

proton mass = 1.6*10^-27 neutron mass = 1.67*10^-27

OpenStudy (fretje):

Beryllium is dangerous. and your data is wrong. 4Be7 is not stable. calculating mass defect for 4Be9: electron mass 9.10938291 × 10-31 kg proton mass 1.67262178 × 10-27 kg neutron mass 1.674894121 × 10-27 kg Be mass 14.96510886 × 10-27 kg mass defect = \[\Delta m\] = 4*1.67262178 × 10^-27 + 5*1.674894121 × 10^-27+4*9.10938291 × 10^-31 - 14.96510886 × 10^-27 = 0.10349261816 × 10^-27

OpenStudy (fretje):

\[\Delta m = 0.10349261816 × 10^{-27} kg\]

OpenStudy (anonymous):

what formula did u use ?

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@nechirwan

OpenStudy (cwrw238):

sorry cany help with this one

OpenStudy (fretje):

so what i did, i added from given data from internet, the mass of all the protons(4 times), all the neutrons(5 times), and all the electrons in the element (4times), and subtract the given mass of the element from the table of Mendeliev. The result is the mass defect. electron mass 9.10938291 × 10-31 kg proton mass 1.67262178 × 10-27 kg neutron mass 1.674894121 × 10-27 kg Beryllium (Be) mass 14.96510886 × 10-27 kg Mass defect = 4*1.67262178 × 10^-27 + 5*1.674894121 × 10^-27+4*9.10938291 × 10^-31 - 14.96510886 × 10^-27 = 0.10349261816 × 10^-27 To understand mass defect one can read http://physics.bu.edu/~duffy/sc546_notes10/mass_defect.html